LUCI1 AO Commissioning

01:43 Opening the dome.

02:29 Working on pointing.

03:39 Seeing still poor, but improving slightly.

04:00 Successful preset bin 3.

Dataset 20160427_040123

04:14 DIMM seeing varying from 1.2" --> 2", but staying below 1.5" mostly. Loop stable.

04:33 Seeing got much worse ~3", stopping AO.

09:54 Beginning AO operations again. Pointing model data were taken with 4hours lost due to telescope colimation problems.

10:15 Bin 2 990Hz, Gopt not working. Data set: 20160427_101528

  • It should be in bin 1, but the WFS estimates the magnitude to be 1 magnitude dimmer than reality.
  • This seems to occur no matter what the magnitude of the star is ie, a 14th mag star is estimated to be 15th magnitude.
  • 110 Modes because I didn't turn off "Bad Seeing Mode"
  • Gopt process was down. Now I think that we would need a new preset in order to send the disturbance again.
10:32 Data Set: 20160427_103128

10:41 Data Set: 20160427_104151

11:20 Wed Apr 27 11:20:18.281 2016 aos.OffsetXY{FLAO}.failed - OffsetXY {FLAO}[151] - Command timeout - -13107: Arbitrator interface generic error

11:30 Wed Apr 27 11:30:02.243 2016 aos.AcquireRef{FLAO}.failed - AcquireRef{FLAO}[157] - Timeout waiting for baysidey.L.ENABLE.CUR (-5001) TIMEOUT_ERROR

-- %USERSIG{GregoryTaylor - 2016-04-27}%
Topic revision: r5 - 27 Apr 2016, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback